An assembly process has a cycle time of 45 seconds and operates with 100 stations.

Assuming the process starts empty and is machine-paced, how long will it take (in minutes) to complete 50 units? (Round your answer to 2 decimal places.)
Answer #1

The first unit will take 45*100 = 4500 seconds and remaining 49 units will have a cycle time of 45 seconds

So, it will take 4500+49*45 = 6705 seconds or 6705/60 = 111.75 minutes to complete 50 units